OpenStudy (anonymous):

How do you use the Factor Theorem to show that the second polynomial is a factor of the first polynomial? 2x^3 - 5x^2 + 6x - 2, x - 1/2

OpenStudy (amistre64):

divide it thru?

OpenStudy (amistre64):

if its got no remainder its a factor

OpenStudy (amistre64):

.5| 2 -5 6 -2 0 1 -2 2 ------------ 2 -4 4 0 <- remainder 0, its a factor

OpenStudy (amistre64):

cant really say that I know what a factor therom is tho ...

OpenStudy (anonymous):

factor theorem sez that if r is a zero of a polynomial \[p(x)\] then \[p(x)=(x-r)q(x)\]

OpenStudy (anonymous):

so if you know \[p(\frac{1}{2})=0\] then you know \[p(x)=(x-\frac{1}{2})q(x)\]

OpenStudy (anonymous):

if f(1/2) is equal to zero then x-1/2 is a factot

OpenStudy (anonymous):

more easily written as \[p(x)=( 2x-1)q(x)\]

OpenStudy (anonymous):

amistre as usual did all the work, so you even know what \[q(x)\] is

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

it is \[2x^2-4x+4=2(x^2+2x+2)\]

OpenStudy (anonymous):

so your polynomial is \[p(x)=(2x-1)(x^2-2x+2)\]

OpenStudy (anonymous):

oops typo in other answer . it is -2x not +2x sorry

OpenStudy (anonymous):

115 to go and then i quit!

OpenStudy (amistre64):

\(\left(\begin{array}ca&b&c\\d&e&f\\g&h&i \end{array} \right)\)

OpenStudy (anonymous):

Thank you